# Tight contact structures without symplectic fillings are everywhere

### 27 Feb 2024 v4

@@ 4.2. Symplectic Cohomology
We will approach Theorem 4.1 through the language of symplectic cohomology. More precisely, for its proof we will need both the existence of certain moduli spaces and the non-existence of others. We then start by recalling the basics of symplectic cohomology (we refer readers to [CO18] for a more detailed account) and describing in detail the moduli spaces involved.
4.2.1. Basics of symplectic cohomology
Let $W$ be a Liouville filling with strict contact boundary $(Y,\alpha)$.
@@ 4.2.1. Basics of symplectic cohomology
    For any $D>0$, which is not the period of some Reeb orbit of $\alpha$, we have a $\mathbb{Z}/2-$graded (with grading given by $\frac{\dim W}{2}-\mu_{CZ}$) filtered positive symplectic cohomology $SH_{+}^{*,<D}(W)$. Its underlying cochain complex is generated by the (hat and check versions of the) Reeb orbits with period up to $D$. We also have the filtered symplectic cohomology $SH^{*,<D}(W)$, whose cochain complex is generated by the above orbits and a Morse cochain complex of $W$. Moreover, these cohomologies fit into a tautological long exact sequence
@@
    We define $\delta_{\partial}:SH_{+}^{*,<D}(W)\to H^{*}(Y)[1]$ to be the composition of the connecting map $\delta$ and the restriction map in cohomology associated to the inclusion $Y\hookrightarrow W$. Then, according to [Zho21b, §3.1], $\delta_{\partial}$ can be defined directly by counting rigid configurations consisting of a holomorphic plane in $W$ and a tail of gradient flow line in $Y$, with respect to any auxiliary Morse function on $Y$. When $\delta_{\partial}(x)$ has a non-trivial component in $H^{d}(Y)$, then we have
@@
    One can also define symplectic cohomology for strong fillings [^2] The existence of positive symplectic cohomology for strong fillings is a consequence of the asymptotic behaviour lemma [CO18, Lemma 2.3]. with coefficient the Novikov field over $\mathbb{Q}$. Regarding the analytical foundation, one needs to apply virtual techniques (e.g. polyfolds, virtual fundamental cycles, Kuranishi techniques) to define symplectic cohomology for general strong fillings going beyond semi-positive symplectic fillings. For our purposes, we can use the construction of Hamiltonian-Floer cohomology for general symplectic manifolds in [Par16], and the results in this paper rely only on the fact that a virtual count/perturbed count can be arranged to be the geometric count when transversality holds, e.g. by [Par19, Proposition 4.33].
4.2.2. Neck-stretching
We will repeatedly apply neck-stretching [${}^{+}$03] to moduli spaces in symplectic cohomology. Roughly speaking, this has two uses:
@@ 4.2.2. Neck-stretching
    exclude certain moduli spaces, by action/index considerations;
  2. (2)
    establish the existence of certain other moduli spaces by comparison with the situation in the “standard filling”.
Both uses will yield a simpler presentation of differentials as well as of other structural maps (e.g. $\delta_{\partial}$) in symplectic cohomology. Here we present a very brief account on how to apply neck-stretching to moduli spaces in symplectic cohomology, and refer readers to [CO18] (see also [Zho21b, Zho22a]) for details.
Let $(Y,\alpha)$ be a contact type hypersurface in a symplectic filling $W$. Assume that $Y$ divides $W$ into the union of a cobordism $X$ from $Y$ to $\partial W$ and a filling $W^{\prime}$ of $Y$. For any almost complex structure $J$, which is compatible with the contact structure near $Y$, we can find a $[0,1)-$family of compatible almost complex structures on $\widehat{W}$ starting at $J_{0}=J$ and converging, for $t\to 1$, to the almost complex structures on the completions $\widehat{X}$, $\widehat{W^{\prime}}$ and $Y\times\mathbb{R}_{+}$ obtained by stretching $J$ in the Liouville vector direction along $Y$. Holomorphic curves and/or Floer cylinders for this family of almost complex structures $J_{t}$, will then converge to SFT buildings for $t\to 1$, in such a way that the top-level curve, i.e. the curve contained in $\widehat{X}$ will develop negative punctures asymptotic to Reeb orbits on $Y$. Let $\Gamma$ be the set of Reeb orbits to which these negative punctures are asymptotic. There are then mainly two constraints on the top-level curve:
  1. (1)
    Action constraint:
    $\mathcal{A}_{H}(\text{positive end})-\mathcal{A}_{H}(\text{negative end})-\sum_{\gamma\in\Gamma}\int\gamma^{*}\alpha>0$
    where $\mathcal{A}_{H}$ is the symplectic action [Zho21b, (2.1)] for the Hamiltonian $H$, which is well-defined if $\widehat{X}$ is exact.
@@
    $m-\sum_{\gamma\in\Gamma}\left(\mu_{CZ}(\gamma)+\frac{\dim W}{2}-3\right)\geq 0,$
@@
    where $m$ is the expected dimension of the moduli space without negative punctures and $\mu_{CZ}(\gamma)+\frac{\dim W}{2}-3$ is the SFT degree of $\gamma$. Note that when we speak of expected dimension we will always consider the unparametrised moduli space where we quotient out any translation symmetries.
We now want to explicitly describe the moduli spaces that we are reduced to count when applying neck-stretching to the contact boundary of a symplectic manifold (recall the discussion in (3) of Section 4.2), seen inside the completion of the latter.
@@
For this purpose, let $Y$ be a contact manifold equipped with a non-degenerate contact form $\alpha$. Let also $H$ be a (time-dependent) Hamiltonian on $\widehat{Y}:=\mathbb{R}_{t}\times Y\simeq(\mathbb{R}_{+})_{r}\times Y$, where $r=e^{t}$, such that $H$ is 00 on $(0,1]_{r}\times Y$ and is the same as the standard Hamiltonian with slope $D$ on the cylindrical end as in the usual definition of symplectic cohomology. The almost complex structure $J$ is also assumed to be cylindrical in the cylindrical end, as in the usual symplectic cohomology setup. We will refer to such a pair $(H,J)$ as admissible.
@@
In this setting, given a multiset $\Gamma=\{\gamma_{1},\ldots,\gamma_{k}\}$ of good Reeb orbits and $x,y$ non-constant Hamiltonian orbits, we can consider the compactification ¯$\overline{{\mathcal{M}}}_{Y,H}(x,y,\Gamma)$ of the moduli space
${\mathcal{M}}_{Y,H}(x,y,\Gamma)={\raisebox{5.0pt}{$\left\{u:\mathbb{R}_{s}\times S^{1}_{t}\backslash\{p_{1},\ldots,p_{k}\}\to\widehat{Y}\;\left|\begin{array}[]{c}\partial_{s}u+J(\partial_{t}u-X_{H})=0,\;\lim_{s\to\infty}u(s,\cdot)=x,\\ \lim_{s\to-\infty}u(s,\cdot)=y,\;\lim_{p_{i}}u=\gamma_{i}\end{array}\right.\right\}$}\left/\raisebox{-5.0pt}{$\mathbb{R}$}\right.}\,.$
Here, the notation $\lim_{p_{i}}u=\gamma_{i}$ means that $u$ is asymptotic to $\gamma_{i}$ at $p_{i}$ viewed as a negative puncture with a free asymptotic marker; recall that in this region the Hamiltonian is zero and the equation is the usual Cauchy-Riemann equation. Note that the compactification ¯$\overline{{\mathcal{M}}}_{Y,H}(x,y,\Gamma)$ is a mixture of Floer-type breaking at non-constant Hamiltonian orbits, as no breaking can happen at a constant orbit of $H$ by symplectic action reasons, and of SFT building breaking at the lower level. This is the moduli space that may appear after applying neck-stretching to the moduli space defining the differential for $SH^{*}_{+}$.
Similarly, for a given singular chain $C\subset Y$ and a fixed $0<\eta\ll 1$, we consider the compactification ¯$\overline{{\mathcal{M}}}_{Y,H}(x,C,\Gamma)$ of the following moduli space:
${\mathcal{M}}_{Y,H}(x,C,\Gamma)={\raisebox{5.0pt}{$\left\{u:\mathbb{C}\backslash\{p_{1},\ldots,p_{k}\}\to\widehat{Y}=\mathbb{R}_{+}\times Y\left|\begin{array}[]{c}\partial_{s}u+J(\partial_{t}u-X_{H})=0,\;\,p_{i}\neq 0,\\ \displaystyle\lim_{s\to\infty}u(s,\cdot)=x,\;\,\lim_{p_{i}}u=\gamma_{i},\\ u(0)\in\{1-\eta\}\times C,\end{array}\right.\right\}$}\left/\raisebox{-5.0pt}{$\mathbb{R}$}\right.}\,,$
In actual constructions, to avoid working with the infinite-dimensional model $C^{*}(Y)$ of singular chains, one can use a finite-dimensional one: for instance, one can just use simplicial complexes from a triangulation or Morse complexes thanks to the choice of an auxiliary Morse function on $Y$, in which case $C$ is replaced by the stable manifolds of the gradient flow. This is the moduli space that may appear after applying neck-stretching to the moduli space defining $\delta_{\partial}$.

### 14 Mar 2026 v5

@@ 4.2. Symplectic Cohomology
As outlined above, the main tool to prove Theorem 4.1 will be symplectic cohomology. Moreover, we will need a precise description of certain moduli spaces. We then start by recalling the basics of symplectic cohomology (we refer readers to [CO18] for a more detailed account).
Basic Properties
Let $W$ be a Liouville filling with strict contact boundary $(Y,\alpha)$ (i.e. we fix the contact form $\alpha$ such that the Liouville form of $W$ is restricted to be $\alpha$ on the boundary).
@@
    (Filtered Symplectic Cohomology) For any $D>0$, which is not the period of some Reeb orbit of $\alpha$, we have a $\mathbb{Z}/2$-graded filtered positive symplectic cohomology $SH_{+}^{*,<D}(W)$. We may assume Reeb orbits of period at most $D$ are non-degenerate. Its underlying cochain complex is generated by the hat $\hat{\gamma}$ and check $\check{\gamma}$ versions of the Reeb orbits with period up to $D$, whose grading is given by $\frac{\dim W}{2}-\mu_{CZ}\mod 2$; here, $\mu_{CZ}(\check{\gamma})=\mu_{CZ}(\hat{\gamma})-1=\mu_{CZ}(\gamma)$ with $\mu_{CZ}(\gamma)$ being the Conley-Zehnder index of the Reeb orbit $\gamma$. The direct limit w.r.t. $D\to+\infty$ yields the symplectic cohomology of $W$, which is an invariant of $W$ up to Liouville homotopy. On the other hand, the filtered positive symplectic cohomology $SH_{+}^{*,<D}(W)$ depends on the boundary contact form $\alpha$ and is not an invariant of $W$ up to homotopy. Note that we suppress the appearance of $\alpha$ into the assumption that $W$ is a Liouville filling with strict contact boundary $(Y,\alpha)$ for simplicity. We also have the filtered symplectic cohomology $SH^{*,<D}(W)$, whose cochain complex is generated by the above orbits, and a Morse cochain complex of $W$. Moreover, these cohomologies fit into a tautological long exact sequence:
    $\ldots\to H^{*}(W)\to SH^{*,<D}(W)\to SH_{+}^{*,<D}(W)\stackrel{{\scriptstyle\delta}}{{\to}}H^{*}(W)[1]\to\ldots.$
@@
    (Boundary Map) We define $\delta_{\partial}:SH_{+}^{*,<D}(W)\to H^{*}(Y)[1]$ to be the composition of the connecting map $\delta$ and the restriction map in cohomology associated to the inclusion $Y\hookrightarrow W$. Then, according to [Zho21b, §3.1], $\delta_{\partial}$ can be defined directly by counting rigid configurations consisting of a holomorphic plane in $W$ and a tail of gradient flow line in $Y$, with respect to any auxiliary Morse function on $Y$. When $\delta_{\partial}(x)$ has a non-trivial component in $H^{d}(Y)$, then we have
@@
    (Novikov Coefficients) Symplectic cohomology for Liouville fillings is classically defined for Liouville fillings over $\mathbb{Z}$, and more generally for semi-positive symplectic fillings with coefficients in the Novikov ring over $\mathbb{Z}$. Appealing to virtual perturbation techniques (e.g. polyfolds, virtual fundamental cycles, Kuranishi techniques) one can furthermore define symplectic cohomology in the case of general symplectic fillings; the drawback is that it is now defined over the Novikov field $\Lambda$ over $\mathbb{Q}$, where
[…]
    This can be done also for positive symplectic cohomology, leveraging the asymptotic behaviour lemma [CO18, Lemma 2.3].
    For the purposes of this paper, it is enough to use the construction of Hamiltonian-Floer cohomology for general symplectic manifolds in [Par16]: more precisely, the results in this paper rely only on the fact that a virtual count/perturbed count can be arranged to be the geometric count when transversality holds, e.g. by [Par19, Proposition 4.33] in the case of Pardon’s VFC. One then also has an analogous tautological exact sequence (with Novikov coefficients):
    $\ldots\to H^{*}(W,\Lambda)\to SH^{*,<D}(W,\Lambda)\to SH_{+}^{*,<D}(W,\Lambda)\stackrel{{\scriptstyle\delta}}{{\to}}H^{*}(W,\Lambda)[1]\to\ldots.$
    The Novikov coefficient cohomology of $W$ is simply the tensor product $H^{*}(W)\otimes_{\mathbb{Q}}\Lambda$.
4.3. Neck-stretching
We will repeatedly apply neck-stretching [BEH+03] to moduli spaces in symplectic cohomology. Roughly speaking, we will use this to reduce the computation of boundary/connecting maps to counting certain curves in the symplectization (with an appropriate choice of Hamiltonian) and hence in a standard model which is “sufficiently stretched”. Here we present a very brief account of how to apply neck-stretching to moduli spaces in symplectic cohomology, and refer readers to [CO18, Lemma 2.4, Proposition 9.17 ] for further details (see also [Zho21b, §3.2] or [Zho23b, §2.5]).
Let $(Y,\alpha)$ be a contact-type hypersurface in a symplectic filling $W$. Assume that $Y$ divides $W$ into the union of a cobordism $X$ from $Y$ to $\partial W$ and a filling $W^{\prime}$ of $Y$.
For any almost complex structure $J$, which is compatible with the contact structure near $Y$, we can find a $[0,1)$-family of compatible almost complex structures on $\widehat{W}$ starting at $J_{0}=J$ and converging, for $t\to 1$, to the almost complex structures on the completions $\widehat{X}$, $\widehat{W^{\prime}}$ and $Y\times\mathbb{R}_{+}$ obtained by stretching $J$ in the Liouville vector direction along $Y$.
$u$
$u$
$u^{\infty}$$\widehat{X}$$\widehat{Y}$$\widehat{W^{\prime}}$
[…]
Holomorphic curves and/or Floer cylinders for this family of almost complex structures $J_{t}$, will then converge to SFT buildings for $t\to 1$, in such a way that the top-level curve, i.e. the curve contained in $\widehat{X}$ will develop negative punctures asymptotic to Reeb orbits on $Y$, see Figure 2. Denoting by $\Gamma$ the set of Reeb orbits to which these negative punctures are asymptotic, we then have the following two constraints on the top-level curve:
@@ 4.3. Neck-stretching
    Action constraint: action increases along the top-level curve, i.e.
    $\mathcal{A}_{H}(\text{output end})-\mathcal{A}_{H}(\text{input end})-\sum_{\gamma\in\Gamma}\int\gamma^{*}\alpha\;\geq 0\,$, (3)
    and the equality holds if and only if:
    • the top-level curve $u_{\infty}$ on $u^{-1}_{\infty}(Y\times(0,1))$ is contained in $\gamma\times(0,1)$, where $Y\times(0,1)_{r}$ is the negative end of the symplectization $Y\times(-\infty,0)_{s}$ with $r=e^{-s}$ and $\gamma$ is a Reeb orbit on $Y$,
    • and $u_{\infty}$ is independent of $s$ on $u_{\infty}^{-1}(\widehat{X}\backslash Y\times(0,1))\subset\mathbb{R}_{s}\times\mathbb{S}^{1}_{t}$.
    In the above inequality, $\mathcal{A}_{H}$ is the symplectic action, which is well-defined if $\widehat{X}$ is exact, for the (time-dependent) Hamiltonian $H$. Here $H$ is zero on $X\cup Y\times(0,1)$ and with a certain slope near the positive cylindrical end of $\widehat{X}$ as in the setup of symplectic cohomology used in [Zho23a]. Namely, if $\lambda$ denotes the Liouville form, for any contractible loop $x\colon\mathbb{S}^{1}\to X$ one has
    $\mathcal{A}_{H}(x)=-\int_{S^{1}}x^{*}\lambda+\int_{S^{1}}H_{t}\circ x(t)\mathrm{d}t\,.$
    As an explicit proof of this inequality is hard to find in the literature, we give the argument below for the reader’s convenience.
@@
    $m-\sum_{\gamma\in\Gamma}\left(\mu_{CZ}(\gamma)+\frac{\dim W}{2}-3\right)\geq 0\,$, (4)
    where $m$ is the expected dimension of the moduli space without negative punctures (which can be expressed using the Conley-Zehnder indices of the asymptotic Hamiltonian orbits depending on the moduli space under consideration), and $\mu_{CZ}(\gamma)+\frac{\dim W}{2}-3$ is the SFT degree of $\gamma$, see e.g. [CO18, Proof of Proposition 9.17].
@@
    Note that when we speak of the expected dimension, we will always consider the unparametrized moduli space where we quotient out any translation symmetries.
    The constraint in (4) follows from a transversality argument, as the top-level curve in all the situations we will consider involves Hamiltonian perturbations and transversality can be achieved for a generic almost complex structure, as those Floer curves have somewhere injective points.
Proof of the action constraint.
In this proof, we write $u$ for the top-level curve $u_{\infty}$ for simplicity. First, notice that there is an inequality
$\int_{u^{-1}(\widehat{X}\backslash Y\times(0,1))}|\partial_{s}u|^{2}+\int_{u^{-1}(Y\times(0,1))}u^{*}{\mathrm{d}}\alpha\geq 0\,$, (5)
where $Y\times(0,1)_{r}$ is the negative end of the completion $\widehat{X}$ with Liouville form $r\alpha$. The non-negativity of the first integral is obvious and the one of the second follows from the fact that we take $H$ to be zero on $Y\times(0,1)$, as in the setup of symplectic cohomology in [Zho21b], so that there $u$ solves the Cauchy-Riemann equation with respect to $J$, that is tamed by ${\mathrm{d}}\alpha$ on $\xi$.
Note that the $s=+\infty$ end of the punctured cylinder in the Floer equation is the input end. Then, because $u$ solves the Floer equation and $\lambda|_{Y\times\{1\}}=\alpha$, using Stokes’ theorem we can rewrite
$\int_{u^{-1}(\widehat{X}\backslash Y\times(0,1))}|\partial_{s}u|^{2}{\mathrm{d}}s{\mathrm{d}}t$ $=$ $\int_{u^{-1}(\widehat{X}\backslash Y\times(0,1))}\omega(\partial_{s}u,J_{t}\partial_{s}u){\mathrm{d}}s{\mathrm{d}}t$
$=$ $\int_{u^{-1}(\widehat{X}\backslash Y\times(0,1))}\omega(\partial_{s}u,\partial_{t}u-X_{H_{t}}(u)){\mathrm{d}}s{\mathrm{d}}t$
$=$ $\int_{u^{-1}(\widehat{X}\backslash Y\times(0,1))}u^{*}{\mathrm{d}}\widehat{\lambda}-{\mathrm{d}}H_{t}(\partial_{s}u){\mathrm{d}}s{\mathrm{d}}t$
$=$ $\mathcal{A}_{H}(\text{output end})-\mathcal{A}_{H}(\text{input end})-\int_{u^{-1}(Y\times\{1\})}\alpha$
The last equality follows from Stokes’ theorem and the fact that $H_{t}$ vanishes on $u^{-1}(Y\times\{1\})$. Here, we implicitly assume $u^{-1}(Y\times\{1\})$ is cut out transversely; if not, by Sard’s theorem, one can just consider $u^{-1}(Y\times\{1-\delta\})$ for a generic $0<\delta\ll 1$ instead.
Similarly, the second integral is
$\int_{u^{-1}(Y\times(0,1))}u^{*}{\mathrm{d}}\alpha=\int_{u^{-1}(Y\times\{1\})}\alpha-\sum_{\gamma\in\Gamma}\int\gamma^{*}\alpha\;.$ (6)
The claimed action constraint then follows directly by adding the above two expressions. The equality case also follows from the fact that the first integrand in (5) is the squared norm of $\partial_{s}u$, and the non-negative form $u^{*}{\mathrm{d}}\alpha$ is 0 if and only if ${\mathrm{d}}u$ is in $\langle\,R_{\alpha},\partial_{r}\,\rangle$, where $R_{\alpha}$ is the Reeb vector field of $\alpha$. ∎
4.4. Model for computing the connecting map $\delta_{\partial}$
We give an explicit description of the moduli spaces that arise when we apply neck-stretching along the contact boundary of a symplectic manifold to the curves used to define the connecting map $\delta_{\partial}$.
@@ 4.4. Model for computing the connecting map $\delta_{\partial}$
Let $Y$ be a contact manifold equipped with a non-degenerate contact form $\alpha$. In addition let $H$ be a (time-dependent) Hamiltonian on $\widehat{Y}:=\mathbb{R}_{t}\times Y\simeq(\mathbb{R}_{+})_{r}\times Y$, where $r=e^{t}$, such that $H$ is 0 on $(0,1]_{r}\times Y$ and is the same as the standard Hamiltonian with slope $D$ on the cylindrical end $2,+\infty)\times Y$ as in the usual definition of symplectic cohomology. The almost complex structure $J$ is also assumed to be cylindrical in the cylindrical end, as in the usual symplectic cohomology setup. We will refer to such a pair $(H,J)$ as admissible.
@@
Given a multiset $\Gamma=\{\gamma_{1},\ldots,\gamma_{k}\}$ of Reeb orbits and non-constant Hamiltonian orbits $x,y$, we consider the moduli spaces
[…]
Here, the notation $\lim_{p_{i}}u=\gamma_{i}$ means that $u$ is asymptotic to $\gamma_{i}$ at $p_{i}$ viewed as a negative puncture with a free asymptotic marker mapped to a chosen base point on $\mathrm{im}\gamma_{i}$; recall that in this region the Hamiltonian is zero and the equation is the usual Cauchy-Riemann equation. The $\mathbb{R}$-action acts on the domain curve, which will move $p_{i}$ as well.
In the cases of interest for us, these moduli spaces will either be smooth 0-dimensional manifolds or will have negative virtual dimension and thus be empty for a generic choice of admissible $(H,J)$. Moreover, in this case, the compactification $\overline{{\mathcal{M}}}_{Y,H}(x,y,\Gamma)$ of ${\mathcal{M}}_{Y,H}(x,y,\Gamma)$ is a mixture of Floer-type breaking at non-constant Hamiltonian orbits into curves of the form (7) and of SFT building breaking at the lower level. Note that no breaking can occur at a constant orbit of $H$ due to the symplectic action constraints. We also point out that SFT-type breaking could appear as follows: if in a sequence of curves the $s$-coordinate of the puncture $p_{i}$ is going to -$\infty$ or $\infty$, then after bringing back $p_{i}$ to a compact part of the cylinder using the $\mathbb{R}$-translation we must see a Floer-type breaking; this particular moduli space of broken configurations is precisely what may appear as the top-level curve after applying neck-stretching to the Floer cylinders corresponding to the differential from $x$ to $y$ in the positive symplectic cohomology.
Similarly, we can consider the neck-stretching limit of curves defining $\delta_{\partial}:SH^{*}_{+}(W)\to H^{*}(\partial W=Y)[1]$. We fix a Morse function $f$ on $Y$ along with a generic metric, for a fixed $\eta>0$ sufficiently small and a critical point $p$ of $f$, we will need to consider the compactification $\overline{{\mathcal{M}}}_{Y,H}(x,S_{p},\Gamma)$ of the following moduli space:
[…]
where $S_{p}$ is the stable manifold of $p$, i.e. points that converge to $p$ under the gradient flow of $f$, and the meaning of $\lim_{p_{i}}u=\gamma_{i}$ is the same as before. In other words, elements in ${\mathcal{M}}_{Y,H}(x,S_{p},\Gamma)$ consist of a curve $u$ from a punctured plane that is asymptotic to the Hamiltonian $x$ at the input end and to Reeb orbits at the extra negative punctures modulo the $\mathbb{R}$-translation along with a tail of gradient line of $f$ from $u(0)$ to $p$ in $\{1-\eta\}\times Y$. The compactification $\overline{{\mathcal{M}}}_{Y,H}(x,S_{p},\Gamma)$ involves Floer breaking at the input end into curves from (7) and (8), SFT building breaking at the negative punctures as well as Morse flow line breaking at the tail. Note that in this setting degenerations where one of the punctures goes to $0\in\mathbb{C}$ can also happen, and they lead to SFT/Floer buildings as discussed in the case of the previous moduli space for a puncture escaping at infinity in the cylinder.
4.5. Reeb dynamics and symplectic cohomology of connected sums
The Reeb dynamics that result from a connected sum can be described explicitly and were worked out by Yau [Yau04], see also [Laz20a]. The following can be proved by combining [Yau04, §4] and [Laz20a, Proof of Proposition 4.4].
Lemma 4.3.
Let $(Y_{1},\alpha_{1})$ and $(Y_{2},\alpha_{2})$ be two $(2n+1)$-dimensional contact manifolds with contact forms whose Reeb orbits do not cover the whole manifold $Y_{1},Y_{2}$. Then, by picking two points $p_{1}\in Y_{1},p_{2}\in Y_{2}$ that do not lie on Reeb orbits, we can attach a Weinstein $1$-handle to them to obtain a contact connected sum $(Y_{1}\#Y_{2},\alpha_{1}\#_{D}\alpha_{2})$ for $D\gg 0$ with the following properties:
  1. (1)
    All Reeb orbits of $\alpha_{1}\#_{D}\alpha_{2}$ of period at most $D$ are either from $Y_{1}$ and $Y_{2}$ outside the handle attachment region or multiple covers $\gamma_{h,i}^{k}$ of the Reeb orbit $\gamma_{h,i}$ for $1\leq i\leq n$ contained in a standard contact sphere of dimension 2n12n-1 in the co-core of the $1$-handle, for $k\leq k_{0}=\lfloor D/\int\gamma_{h,i}^{*}(\alpha_{1}\#_{D}\alpha_{2})\rfloor$ (for any $i$, as $\gamma_{h,i}$ has approximately the same small period).
  2. (2)
    The Conley-Zehnder index of $\gamma_{h,i}^{k}$ is $n-1+2(k-1)n+2i\geq n+1$, where the Conley-Zehnder index is computed from bounding discs contained in the handle. That is, the Conley-Zehnder index of those orbits enumerates through $\{n+1,n+3,\ldots,2k_{0}n+n-1\}$. It is the canonical global Conley-Zehnder index if $c_{1}(Y_{1})=c_{1}(Y_{2})=0$.
The handle grows thinner as $D\to+\infty$.
The purpose of assuming that the base points $p_{1},p_{2}$ are not on a closed Reeb orbit is to find sufficiently small Darboux neighborhoods such that the Reeb flow from the Darboux ball will not return within time $D$. Those orbits $\gamma_{h,i}^{k}$ are from a standard contact sphere of codimension $2$ contained in the co-core of the $1$-handle with an ellipsoid contact form that is close to a round contact form; hence there are precisely $n$ simple Reeb orbits with approximately the same period. Their Conley-Zehnder index is precisely the Conley-Zehnder index on the ellipsoid of dimension 2n12n-1, as the linearized flow in the symplectic normal direction is positive hyperbolic, see [Yau04, Lemma 3.1, Theorem 3.1] for the computation in the handle model.
The effect of contact connected sum on symplectic cohomology is well-understood by the work of Cieliebak [Cie02b]. The following is just a filtered version of the corresponding result in symplectic cohomology and was implicitly proved in [Cie02b]; see [CO18, Proposition 9.19] for an equivalent statement formulated using symplectic cohomology of cobordisms. We only state a special case for the purpose of Theorem 4.1.